OpenStudy (anonymous):

Can someone help me? Rewrite the expressions using a radical. 25x^3/2 and (25x)^3/2

OpenStudy (mertsj):

Remember: the bottom number is the root. The top number is the power.

OpenStudy (mertsj):

\[25x ^{\frac{3}{2}}=25\sqrt{x ^{3}}=25x \sqrt{x}\]

OpenStudy (anonymous):

hey thx i got both questions now :)

OpenStudy (mertsj):

\[(25x)^{\frac{3}{2}}=\sqrt{(25x)^3}=\sqrt{25\times 25\times 25\times x ^{2}\times x}=\]

OpenStudy (mertsj):

\[25x \sqrt{25x}\]

OpenStudy (mertsj):

yw